# Geometric Bounds for Persistence

### v1

@@ Abstract
In this paper, we bring a new perspective to persistent homology by incorporating key concepts from metric geometry. For a given compact subset $\mathcal{X}$ of a Banach space $\mathbf{Y}$, we study the topological features appearing in family $\mathcal{N}_{\bullet}(\mathcal{X}\subset\mathbf{Y})$ of nested of neighborhoods of $\mathcal{X}$ in $\mathbf{Y}$ and give several geometric bounds on their persistence (lifespans).
First, we discuss the lifespans of these homology classes in terms of their filling radii in $\mathbf{Y}$. By using this, we relate these lifespans to key invariants in metric geometry such as Urysohn width. Next, we bound these lifespans via $\ell^{\infty}$-principal components of the set $\mathcal{X}$, which are also known as Kolmogorov widths. Furthermore, we introduce and study the notion of extinction time of a space $\mathcal{X}$: the critical threshold after which there are no homological features alive in any degree. We also describe novel approaches on how to estimate Čech and Vietoris–Rips extinction times of a set $\mathcal{X}$ by relating $\mathcal{X}$ to its über-contractible cores and to its tight span, respectively.
@@ 1 Introduction
In the past decade, several approaches within Topological Data Analysis (TDA) have been developed in order to facilitate the discovery of patterns in various forms of data. Persistent Homology (PH) – one of the the key concepts in TDA – provides an effective multiscale feature extraction method. In particular, PH has become quite popular in various machine learning tasks, ranging from bioinformatics to finance, materials science to network analysis  [${}^{+}$18, HMR21, ${}^{+}$20, SL22, Giu22, PLX22].
In this paper, we study some aspects of the PH methodology with tools from metric geometry. Our results provide novel quantitative interpretations of the PH output.
@@
In many applications, the lifespan of a topological feature $\omega$ has crucial significance as it is considered to reflect the size or degree of “importance” of $\omega$. Indeed, in practice, the topological features with long lifespans are thought to represent the main shape characteristics of a dataset whereas the ones with short lifespans are usually regarded as (topological) noise. Therefore, determining and interpreting the lifespans of these topological features appearing in the persistence diagram has critical significance for applications.
In this paper, we aim to give a geometric interpretation of these lifespans by relating them to several notions from metric geometry. We study the lifespans of individual homology classes $\omega$ appearing throughout the Vietoris-Rips and Čech filtrations by resorting to the notion of filling radius and to several notions of width. Our results show that the lifespans of these homology classes are determined by the geometric size of their representatives in the filtration. We first discuss implications of the the absolute (Gromov’s) and relative filling radius of a homology class (2.30) in our setting. Then, we observe that the Čech lifespan of a homology class $\omega$ is equal to its relative filling radius in its ambient space (Section 2.3).  Next, we give several bounds for the lifespans of individual homology classes by resorting to the notions of Urysohn width, Alexandrov width, and Kolmogorov width. Since all these notions of width are monotonically non-increasing with respect to the dimension parameter (e.g., $\operatorname{UW}_{k}(\mathcal{X})\geq\operatorname{UW}_{k+1}(\mathcal{X})$), any estimate for degree-$k$ automatically applies to homology classes in higher degrees.
@@
$d_{\omega}-b_{\omega}\leq\operatorname{UW}_{k-1}(\overline{\mathcal{N}}_{b_{\omega}}(\mathcal{X}\subset\operatorname{E}(\mathcal{X}))).$
In particular, $d_{\omega}-b_{\omega}\leq\operatorname{UW}_{k-1}(\operatorname{E}(\mathcal{X}))$.
Here, $\mathcal{N}_{r}(\mathcal{X}\subset\mathbf{Y})$ denotes the open $r$-neighborhood of $\mathcal{X}$ in $\mathbf{Y}$ and $\operatorname{E}(\mathcal{X})$ denotes the tight span of $\mathcal{X}$. Notice that in a special case, if $\mathcal{X}$ is a closed $k$-manifold, and $\omega=[\mathcal{X}]$ is the the fundamental class corresponding of $\mathcal{X}$, since $b_{\omega}=0$ in that case, the result above implies that Vietoris-Rips lifespan of $\omega$ is bounded above by the Urysohn width of $\mathcal{X}$, i.e., $d_{\omega}\leq\operatorname{UW}_{k-1}(\mathcal{X})$ (this particular bound goes back to Gromov; it follows from [Gro83, Appendix 1, Example after Lemma (B)] combined with [Gro83, Appendix 1, Proposition (D)]).
@@
In particular, $d_{\omega}-b_{\omega}\leq\operatorname{AW}_{k-1}(\overline{\mathrm{conv}}(\mathcal{X})\subset\mathbf{Y})$.
@@
Further, we introduce a new notion of width, called treewidth (3.9), in order to obtain a finer estimate on Čech lifespans (3.12). Notice that in the estimates above, for each class $\omega$, in order to get an upper bound for its lifespan, the corresponding width needs to be calculated for $\mathcal{N}_{b_{\omega}}(\mathcal{X}\subset\cdot)$, not $\mathcal{X}$ itself. In the following, we get rid of this dependency. The notion of treewidth permits establishing a certain multiplicative bound on lifespans.
@@
[Čech Lifespans via Treewidth]corcortw Let $\mathcal{X}$ be compact subset in a Banach space $\mathbf{Y}$ and let $\omega\in\mathrm{Spec}_{k}(\mathcal{C}_{\bullet}(\mathcal{X}\subset\mathbf{Y}))$, with the birth time $b_{\omega}\geq 1$. Then,
@@
Next, we consider the $\ell^{\infty}$-version of principal component analysis (${}_{\infty}$) for a compact subset $\mathcal{X}$ of a Banach space $\mathbf{Y}$ (e.g., a point cloud in $\mathbb{R}^{N}$). By using the estimates in previous sections, we relate the $(k+1)$-variance $\nu_{k+1}(\mathcal{X}\subset\mathbf{Y})$ with the lifespans of classes $\omega$ appearing throughout the Čech filtration (Section 3.5). We note that the variance $\nu_{k+1}(\mathcal{X}\subset\mathbf{Y})$ is also known as the $k^{th}$ Kolmogorov width of $\mathcal{X}$ in the approximation theory literature (3.24).
@@
[Čech lifespans via $\infty$-Variance]corcorpca Let $\mathcal{X}$ be a compact subset of a Banach space $\mathbf{Y}$ and let $\omega\in\mathrm{Spec}_{k}(\mathcal{C}_{\bullet}(\mathcal{X}\subset\mathbf{Y}))$, $k\geq 0$. Then,
@@
Note that by monotonicity of widths (e.g., $\operatorname{AW}_{k}(\mathcal{X}\subset\mathbf{Y})\geq\operatorname{AW}_{k+1}(\mathcal{X}\subset\mathbf{Y}))$, all these bounds apply to homology classes of degree higher than $k$, too. While these lifespan bounds depend on the homology degree, we next give a general bound for Čech lifespans which is independent of degree. We achieve this by generalizing Katz’s notion of spread to the extrinsic setting and introducing the notion that we call überspread (Section 4). Überspread basically measures the Hausdorff distance of the space to the closest übercontractible space (a contractible space where all neighborhoods are also contractible) in an ambient space. With this notion, we generalize the existing VR-lifespan estimates with spread in [LMO23] to Čech-lifespans in any degree.
@@
[Čech Lifespans via Überspread]thmthmuberspread Let $\mathcal{X}$ be a compact subset of a Banach space $(\mathbf{Y},\|\cdot\|)$. Let $\omega\in\mathrm{Spec}_{k}(\mathcal{C}_{\bullet}(\mathcal{X}\subset\mathbf{Y}))$ for any $k\geq 0$. Then,
@@
[Bounding Čech Extinction]thmthmcechextintion Let $\mathcal{X}$ be a compact subset of a Banach space $(\mathbf{Y},\|\cdot\|)$. Then,
@@
From a conceptual perspective, our results can be considered as relating a given space $\mathcal{X}$ with another space $\Lambda_{\mathcal{X}}$ (its core). Estimates on the lifespans of homological features created in the course of increasing the radius of neighborhoods of $\mathcal{X}$ are obtained through the distance between $\mathcal{X}$ and $\Lambda_{\mathcal{X}}$. In our study of different notions of widths (Section 3), the role of core $\Lambda_{\mathcal{X}}$ is, roughly speaking, played by the $k$-dimensional space which is closest to $\mathcal{X}$ and $k$-width can be regarded as the distance between $\mathcal{X}$ and $\Lambda_{\mathcal{X}}$.In the section where we consider the notion of spread (Section 4), the core $\Lambda_{\mathcal{X}}$ becomes an übercontractible space and the distance between $\mathcal{X}$ and $\Lambda_{\mathcal{X}}$ upper bounds the lifespans in any degree. Here, the condition on the core is having trivial homology groups rather than a condition on its dimension. Similarly, in the section where we consider extinction times (Section 5), we enforce strong geometric conditions (convexity or hyperconvexity) on $\Lambda_{\mathcal{X}}$ again to obtain our extinction bounds through the distance between $\Lambda_{\mathcal{X}}$ and $\mathcal{X}$.
We would like to point out that from one side, while we give upper bounds for quantities from applied topology (e.g., lifespans of homology classes) with metric geometry concepts, at the same time, these results provide computational lower bounds for metric geometry notions based on ideas from persistent homology. This bidirectional relationship underscores the synergy between these fields, enabling us to derive insights that contribute to both domains.
@@ 1.1 Related Work
In this work, we aim to build a bridge between two seemingly far fields, i.e., applied algebraic topology and metric geometry. In a way, both applied algebraic topology and metric geometry are interested in similar question with different tools: How to measure the size of a set/space/manifold? In applied algebraic topology, persistent homology is a very effective tool for accomplishing this aim, and the lifespans (or persistence) of topological features induced by Vietoris–Rips (or Čech) filtrations represent the size of the corresponding topological features. In particular, Vietoris–Rips (or Čech) complexes were invented in order to transform a given metric space into a simplicial complex by maintaining its topological information, thus enabling an effective cohomology theory for metric spaces [Vie27, Hau95, Bor48]. There are several works in the literature which study Vietoris–Rips filtrations in different settings [${}^{+}$09, AA17, AAR19, ${}^{+}$20, AMMW22, CDSO14, LMO23, Vir20, Vir22].
@@
On the other hand, from the metric geometry side, estimating the size of a manifold has been a key problem for years. Gromov introduced and studied the notion of filling radius in his seminal paper [Gro83], and several other notions of “largeness” in [Gro86]. Before Gromov, certain relative, or extrinsic, versions of the filling radius were studied in geometric analysis in connection with the isoperimetric inequality [BS83, FF60, MS73]. Gromov brought the filling radius to the realm of systolic geometry, and the study of scalar curvature [GL83]. Several other measures of size of a given manifold or metric space, now known as widths, were also studied and popularized by Gromov [Gro83, Gro88]. The interplay between filling radii, widths, and other metric invariants (from volume and beyond) has been an active research area since then. Katz determined the filling radius of spheres and other essential spaces [Kat83, Kat89, Kat90, Kat91]. Several authors have studied the filling radii in comparison with other largeness properties [Cai94, BH10]. Guth proved some related conjectures of Gromov [Gut11, Gut17]. Sabourau, Nabutovsky, and Rotman related the filling radius with sweepouts of manifolds [Sab20, NRS21]. Bounds on filling radius in terms of Hausdorff content follow from very general isoperimetric estimates due to Liokumovich, Lishak, Nabutovsky, and Rotman [LLNR22].
@@
In the past couple of years, there have been some articles exploring connections between these two worlds working on analogous problems with different tools. With this aim, Lim, Mémoli, and Okutan related the filling radius of a closed manifold to the bar corresponding to fundamental class in the top VR persistence diagram [LMO23]. As shown in [LMO23, Section 9.3.2], the stability of persistence diagrams of Vietoris-Rips filtrations can be used to obtain lower bounds for the Gromov-Hausdorff distance between spheres which arise from considerations related to their filling radii, which in [LMS22] were shown to not be tight. The polymath project  [${}^{+}$22] recently furthered this line work. Recently, Adams–Coskunuzer used a well-known quantity in metric geometry, Urysohn width, to estimate the lifespans in the persistence diagram of a given space [AC22]. In [Vir22], Virk studied the relation between higher persistence diagrams for large degrees and lower dimensional features in the manifold setting.
@@
In this paper, we aim to establish a direct connection between notions from applied algebraic topology and notions in metric geometry by relating several quantities used to measure the size of a manifold in both subfields.
@@
This work was supported by the following grants from the National Science Foundation (Grants # DMS-1926686, DMS-2202584, DMS-2220613, DMS-2229417, DMS-1723003, IIS-1901360, CCF-1740761, and DMS-1547357), from the Simons Foundation (Grant # 579977) and from the BSF (under grant # 2020124). The project has received funding from the HORIZON EUROPE Research and Innovation programme under Marie Skłodowska-Curie grant agreement number 101107896. A major part of this work was done while the first-named author was a postdoc at the Institute for Advanced Study, Princeton.

### v2

@@ Abstract
In this paper, we offer a new perspective on persistent homology by integrating key concepts from metric geometry. For a given compact subset $\mathcal{X}$ of a Banach space $\mathbf{Y}$, we analyze the topological features arising in the family $\mathcal{N}_{\bullet}(\mathcal{X}\subset\mathbf{Y})$ of nested neighborhoods of $\mathcal{X}$ in $\mathbf{Y}$ and provide several geometric bounds on their persistence (lifespans).
We begin by examining the lifespans of these homology classes in terms of their filling radii in $\mathbf{Y}$, establishing connections between these lifespans and fundamental invariants in metric geometry, such as the Urysohn width. We then derive bounds on these lifespans by considering the $\ell^{\infty}$-principal components of $\mathcal{X}$, also known as Kolmogorov widths.
Additionally, we introduce and investigate the concept of extinction time of a metric space $\mathcal{X}$: the critical threshold beyond which no homological features persist in any degree. We propose methods for estimating the Čech and Vietoris–Rips extinction times of $\mathcal{X}$ by relating $\mathcal{X}$ to its convex hull and to its tight span, respectively.
@@ 1 Introduction
Over the past decade, numerous approaches within Topological Data Analysis (TDA) have been developed to uncover patterns across a wide variety of data types. Among these, Persistent Homology (PH) has emerged as a cornerstone of TDA, providing a robust multiscale feature extraction framework. This progress has been driven by the development of efficient algorithmic procedures and effective software implementations for its computation (see Section 2.1). Indeed, PH has gained considerable traction in diverse machine learning applications spanning fields such as:
In this paper, we explore aspects of the Persistent Homology (PH) methodology through the lens of metric geometry. Our results offer new quantitative interpretations of the PH output, providing deeper insights into its structure and significance.
1.1 Persistent Homology (in a nutshell)
@@
The notion of Persistent Homology is closely related to but subtly differs from this process of recording birth and death times for individual homology classes described above. The degree-$k$ persistent homology of $\Delta_{\bullet}$ is the directed system of vector spaces $\widetilde{\mathrm{H}}_{k}(\Delta_{\bullet};\mathbb{F})$. Under suitable tameness assumptions on the family $\Delta_{\bullet}$, an up-to-isomorphism representation of this directed system can be obtain via its persistence diagram, a multiset of intervals $I$ on $\mathbb{R}_{>0}$ supporting certain linearly independent collection of nontrivial homology classes that are alive at all points in $I$. See Section 2.1 for the precise definition of persistence diagrams (PD).
In many applications, the lifespan of a topological feature $\omega$ is critically significant, as it is often interpreted as a measure of the “size” or “importance” of $\omega$. In practice, topological features with long lifespans—those that persist—are typically considered to represent the primary shape characteristics of a dataset, while features with short lifespans are generally regarded as (topological) noise. [^3] However, short lived, or even ephemeral, topological features also can carry useful information; see Usher and Zhang [UZ16], Bubenik, Hull, Patel and Whittle [BHPW20], and Mémoli and Zhou [MZ24]. Therefore, determining (or estimating) and interpreting the lifespans of these topological features appearing in the persistence diagram has critical significance for applications such as the ones mentioned above.
1.2 Connections with Metric Geometry and Main Results
In this paper, we aim to give a geometric interpretation of these lifespans by relating them to several notions from Metric Geometry. We study the lifespans of individual homology classes $\omega$ appearing throughout the Vietoris–Rips and Čech filtrations by resorting to the notion of filling radius and to several notions of width. See Section 1.3 for a discussion of the interplay between widths and filling radii in metric geometry. Through these notions of width, our results show that the lifespans of homology classes are controlled by the (geometric) size of their representatives in the filtration thus providing precise (geometric) interpretations of the significance of the features tracked by persistent homology.
We first discuss implications of the absolute (Gromov’s) and relative filling radius of a homology class (2.31) in our setting. Then, we observe that the Čech lifespan of a homology class $\omega$ is equal to its relative filling radius in ambient space (Section 2.3).  Next, we give several bounds for the lifespans of individual homology classes by resorting to the notions of Urysohn width, Alexandrov width, and Kolmogorov width. They measure in various ways how well a space can be approximated by a $k$-dimensional complex, and they are denoted by $\operatorname{UW}_{k}(\cdot),\operatorname{AW}_{k}(\cdot),\operatorname{KW}_{k}(\cdot)$, respectively. Since all these notions of width are monotonically non-increasing with respect to the dimension parameter (e.g., $\operatorname{UW}_{k}(\mathcal{X})\geq\operatorname{UW}_{k+1}(\mathcal{X})$), any degree-$k$ estimate automatically applies to homology classes in higher degrees.
@@ 1.2 Connections with Metric Geometry and Main Results
$d_{\omega}-b_{\omega}\leq\operatorname{UW}_{k-1}\big{(}\overline{\mathcal{N}}_{b_{\omega}}(\mathcal{X}\subset\operatorname{E}(\mathcal{X}))\big{)}.$
In particular,
$d_{\omega}-b_{\omega}\leq\operatorname{UW}_{k-1}(\operatorname{E}(\mathcal{X})).$
Here and throughout the paper, $\overline{\mathcal{N}}_{r}(\mathcal{X}\subset\mathcal{Z})$ denotes the closed $r$-neighborhood of $\mathcal{X}$ inside the metric space $\mathcal{Z}$ and $\operatorname{E}(\mathcal{X})$ denotes the tight span of $\mathcal{X}$ (2.28), a canonically constructed metric space admitting an isometric embedding of $\mathcal{X}$ and enjoying properties reminiscent of (but stronger than) the ones of the convex hull.
Notice that in a special case, if $\mathcal{X}$ is a closed $k$-manifold, and $\omega=[\mathcal{X}]$ is its fundamental class, since $b_{\omega}=0$ in that case, the result above implies that the Vietoris–Rips lifespan of $\omega$ is bounded above by the Urysohn width of $\mathcal{X}$, i.e., $d_{\omega}\leq\operatorname{UW}_{k-1}(\mathcal{X})$ (this particular bound goes back to Gromov; it follows from [Gro83, Appendix 1, Example after Lemma (B)] combined with [Gro83, Appendix 1, Proposition (D)]).
@@
In particular,
$d_{\omega}-b_{\omega}\leq\operatorname{AW}_{k-1}(\overline{\mathrm{conv}}(\mathcal{X})\subset\mathbf{Y}).$
@@
Further, we introduce a new notion of width, called treewidth (3.9), in order to obtain a finer estimate on Čech lifespans (3.12). Notice that in the estimates above, for each class $\omega$, in order to obtain an upper bound for its lifespan, the corresponding width needs to be calculated for the neighborhood $\mathcal{N}_{b_{\omega}}(\mathcal{X}\subset\cdot)$ of $\mathcal{X}$, not for $\mathcal{X}$ itself. In the following, we get rid of this dependency on the neighborhood.
The notion of treewidth permits establishing a certain multiplicative bound on lifespans.
@@
[Čech Lifespans via Treewidth]corcortw Let $\mathcal{X}$ be compact subset in a Banach space $\mathbf{Y}$ and let $\omega\in\mathrm{Spec}_{k}(\mathcal{C}_{\bullet}(\mathcal{X}\subset\mathbf{Y}))$, with birth time $b_{\omega}\geq 1$. Then,
@@
[…]
@@
[Čech Lifespans via $\ell^{\infty}$-Variance]corcorpca Let $\mathcal{X}$ be a compact subset of a Banach space $\mathbf{Y}$ and let $\omega\in\mathrm{Spec}_{k}(\mathcal{C}_{\bullet}(\mathcal{X}\subset\mathbf{Y}))$, $k\geq 0$. Then,
@@
Note that by monotonicity of widths (e.g., $\operatorname{AW}_{k}(\mathcal{X}\subset\mathbf{Y})\geq\operatorname{AW}_{k+1}(\mathcal{X}\subset\mathbf{Y}))$, all these bounds apply to homology classes of degree higher than $k$, too.
While these lifespan bounds depend on the homology degree, we next give a general bound for Čech lifespans which is independent of degree. We achieve this by generalizing Katz’s notion of spread [Kat83] to the extrinsic setting and introducing the notion that we call überspread (Section 4). Überspread basically measures the Hausdorff distance from the space to the closest übercontractible space (a contractible space where all neighborhoods are also contractible, see 4.2) in an ambient space. With this notion, we generalize the existing VR-lifespan estimates via spread obtained by Lim, Mémoli and Okutan [LMO24] to Čech-lifespans in any degree.
@@
[Čech Lifespans via Überspread]thmthmuberspread Let $\mathcal{X}$ be a compact subset of a Banach space $\mathbf{Y}$. Let $\omega\in\mathrm{Spec}_{k}(\mathcal{C}_{\bullet}(\mathcal{X}\subset\mathbf{Y}))$ for any $k\geq 0$. Then,
@@
[Bounding Čech Extinction]thmthmcechextintion Let $\mathcal{X}$ be a compact subset of a Banach space $\mathbf{Y}$. Then,
@@
We highlight the bidirectional relationship between applied topology and metric geometry: on one hand, with the goal of improving their interpretability, we establish upper bounds for crucial quantities in applied topology (e.g., lifespans of homology classes) using concepts from metric geometry; on the other hand, these results yield computational lower bounds or estimates for metric geometry notions inspired by persistent homology; see 2.36. This interplay underscores the synergy between these fields, enabling insights that advance both domains.
Cores. Conceptually, our results establish a relationship between a given space $\mathcal{X}$ and another space $\Lambda_{\mathcal{X}}$, which functions as a core for $\mathcal{X}$. This represents the central thread weaving together the various parts of the paper. Specifically, estimates on the lifespans of homological features that arise as the radius of neighborhoods of $\mathcal{X}$ increases are derived from the distance between $\mathcal{X}$ and $\Lambda_{\mathcal{X}}$:
  • In Section 3, where we explore various notions of widths, the role of the core $\Lambda_{\mathcal{X}}$ is, roughly speaking, assumed by a $k$-dimensional space closest to $\mathcal{X}$. In this context, the $k$-width can be interpreted as the distance between $\mathcal{X}$ and this approximate core, $\Lambda_{\mathcal{X}}$. In this section, we first recall the classical notions of Urysohn and Alexandrov width and then introduce a new variant which we call treewidth.
  • In Section 4, after recalling the notion of spread, we introduce the notion of überspread, where the core $\Lambda_{\mathcal{X}}$ is treated as an übercontractible space, and the distance between $\mathcal{X}$ and $\Lambda_{\mathcal{X}}$ provides an upper bound on the lifespans in any degree. Notably, the condition imposed on the core in this context is the triviality of its homology groups, rather than any restriction on its dimension.
  • In Section 5, where we analyze extinction times, we impose strong geometric conditions—such as convexity or hyperconvexity—on $\Lambda_{\mathcal{X}}$ to derive extinction bounds based on the distance between $\Lambda_{\mathcal{X}}$ and $\mathcal{X}$. This is done through the concepts of convexity deficiency and hyperconvexity deficiency, which we introduce therein.
@@ 1.3 Related Work
In this work, we aim to build a bridge between two seemingly disparate fields: applied algebraic topology and metric geometry. Both disciplines address a similar fundamental question regarding the quantification of “shape” using distinct tools:
How to measure the size of a set/space/manifold?
@@
In applied algebraic topology, persistent homology is a very effective tool for accomplishing this aim, and the lifespans (or persistence) of topological features induced by Vietoris–Rips (or Čech) filtrations are used as a measure of the size or importance of the corresponding topological features. In particular, Vietoris–Rips (or Čech) complexes were invented in order to transform a given metric space into a simplicial complex by maintaining its topological information, thus enabling an effective cohomology theory for metric spaces; see the papers by Vietoris [Vie27], Borsuk [Bor48] and Hausmann [Hau95]. Numerous studies in the literature explore Vietoris–Rips complexes and Vietoris–Rips filtrations across various settings; see Latschev [Lat01], Chazal, Cohen-Steiner, de Silva, Guibas, Mémoli, and Oudot [CCSG+09, CDSO14], Adamaszek, Adams, Frick, Gillespie, Lim, Mémoli, Moy, Okutan, Reddy, and Wang [AA17, AAR19, AMMW24, AFV23, LMO24, Gil24], Attali, Lieutier and Salinas [ALS13], Rieser, Bubenik and Milicevic [Rie20, BM24], Turner [Tur19], Virk [Vir20, Vir22, Vir21], and Zaremsky [Zar22].
@@
On the other hand, from the metric geometry side, estimating the size of a manifold has been a key problem for several decades. Gromov introduced and studied the notion of filling radius in his seminal paper [Gro83], and several other notions of “largeness” in [Gro86]. Before Gromov, certain relative, or extrinsic, versions of the filling radius were studied by Federer and Fleming [FF60], Michael and Simon [MS73] and Bombieri and Simon [BS83] in geometric analysis in connection with the isoperimetric inequality. Gromov and Lawson brought the filling radius to the realm of systolic geometry, and the study of scalar curvature [GL83]. Several other measures of size of a given manifold or metric space, nowadays known as widths, were also studied and popularized by Gromov [Gro83, Gro88].
The interplay between filling radii, widths, and other metric invariants (from volume and beyond) has been an active research area since then. Katz determined the filling radius of spheres and other essential spaces [Kat83, Kat89, Kat90, Kat91]. Several authors have studied the filling radii in comparison with other measures of largeness [Cai94, BH10]. Guth proved some related conjectures of Gromov [Gut11, Gut17]. Sabourau, Nabutovsky, and Rotman related the filling radius with sweepouts of manifolds [Sab20, NRS21]. Bounds on filling radius in terms of Hausdorff content follow from very general isoperimetric estimates due to Liokumovich, Lishak, Nabutovsky, and Rotman [LLNR22].
@@
In recent years, several articles have explored connections between these two domains, addressing analogous problems with different methodologies. With this aim, Lim, Mémoli and Okutan related the filling radius of a closed manifold to the bar corresponding to the fundamental class in the top VR persistence diagram [LMO24]. As shown in [LMO24, Section 9.3.2], the stability of persistence diagrams of Vietoris–Rips filtrations can be used to obtain lower bounds for the Gromov–Hausdorff distance between spheres through considerations related to their filling radii. These were shown to not be tight by Lim, Mémoli and Smith [LMS23] and the polymath project  [ABC+22] furthered this line work; see also Jeffs and Harrison [HJ23] and Rodriguez-Martín [RM24]. Recently, Adams and Coskunuzer used a well-known quantity in metric geometry, Urysohn width, to estimate the lifespans in the persistence diagram of a given space [AC22]. In [Vir22], in the manifold setting, Virk studied the relation between persistence diagrams for large degrees and lower dimensional features.
@@
In this paper, we aim to establish a direct connection between concepts from applied algebraic topology and metric geometry by linking various quantities used in both fields to measure the size of a metric space. In writing this paper, we have prioritized accessibility, aiming to bridge the gap between the two fields and foster greater collaboration and understanding.
@@
This work was supported by the following grants from the National Science Foundation (Grants # DMS-1926686, DMS-2202584, DMS-2220613, DMS-2229417, DMS-1723003, IIS-1901360, CCF-1740761, and DMS-1547357), from the Simons Foundation (Grant # 579977) and from the BSF (under grant # 2020124). The project has received funding from the HORIZON EUROPE Research and Innovation programme under Marie Skłodowska-Curie grant agreement number 101107896. A major part of this work was done while the first-named author was a postdoctoral member at the Institute for Advanced Study, Princeton. We thank the anonymous reviewers for their feedback which helped enhance the presentation of this work.